# Black Hole Superradiance of Interacting Multi-Field

### v2

@@ Abstract
We investigate black hole superradiance evolution of the interacting multiple fields. We consider a model of two scalar fields interacting with a cubic coupling, and study the superradiant evolution of the cloud. We demonstrate that superradiance is typically suppressed when the superradiant field couples to another field, even with a very weak coupling strength. This implies that the constraints on dark particles derived from single-field analyses can be revised in the presence of interactions. Moreover, we find that the multi-field superradiant evolution and its corresponding observational signatures can be different across parameter spaces, which makes black hole superradiance an even more powerful probe of the dark sector in particle physics.
@@ I Introduction
While the dark sector may consist of multiple particle species with non-trivial interactions, most studies assume the presence of only a single superradiant field. Although some works have considered interactions between the superradiant field and other fields, their focus has largely been on phenomena, such as photon polarization by the cloud [48, 49, 50] and annihilation of the cloud into photons [51], fermions [52] and scalars [53], with the assumption of a preformed superradiant cloud. Not only has the superradiance of the coupled fields (or generally, interactions between the black hole and bound states of the coupled fields) been overlooked, but its consequential effects on the superradiance of the primary field also lack investigation. Therefore, a systematic investigation of superradiant evolution of the interacting multiple fields is essential and necessary.
In this work, we investigate black hole superradiance in the presence of multiple fields. We consider a model of two scalar fields interacting with a cubic coupling, and study the superradiance evolution of the cloud. We demonstrate that superradiance is typically suppressed if the superradiant field couples to another field even with a very weak coupling strength. This implies that the constraints on dark particles derived from single-field analyses can be significantly revised in the presence of interactions. Moreover, the superradiant evolution in different regions of the parameter space are distinct. This makes black-hole superradiance an even more powerful probe of the dark sector in particle physics.
The rest of this paper is organized as follows. We start with the superradiance evolution of two non-interacting fields in Sec. II. Then we shall consider a cubic interaction, and shall discuss its effects on the superradiant growth rate in Sec. III and the superradiant evolution in Sec. IV. Finally, we shall discussion the implications observations in Sec. V. Sec. VI devotes to conclusion and discussion. We will take (,+,+,+)(-,+,+,+) metric convention and set $\hbar=c=1$.
II Superradiance of free multi-fields
In this section, we investigate the superradiance process of non-interacting multi-fields (also see Ref. [54] for a similar study). For the purpose of demonstration, we shall start with a model of two massive scalar fields in the Kerr background, the Lagrangian of which is given by
@@ II.1 Eigenstates of free fields
where $x^{\mu}=(t,r,\theta,\varphi)$ are the Boyer-Lindquist coordinates, $R_{\omega\ell m}(r)$ is the radial function, and $S_{\omega\ell m}(\theta)$ turns out to be the spheroidal harmonics. The boundary conditions at the black hole horizon and at spatial infinity single out a set of eigenfrequencies. For bound states, i.e., states vanishes at spatial infinity, the eigenfrequencies $\omega$ are shown to be a set of discrete complex number $\omega_{n\ell m}$, labelled by three quantum numbers $n$, $\ell$ and $m$. These eigenfrequencies can be obtained numerically. For instance, Fig. 1 shows the eigenfrequencies of some bound states obtained with the continued fraction method [55]. For $\alpha\ll 1$, we have [14]
@@
Figure 1: Eigenfrequencies of bound states with $n\ell m$ being 211211 (in black) and 322322 (in red). The upper and lower plots show the real and imaginary parts of the eigenfrequencies respectively. The solid lines are given by Eqs. (3) and (4), while the dashed lines show the resultes obtained from numerical calculation.
While the wavefunction of the bound states can also be obtained numerically [62], for $\alpha\ll 1$, it is convenient to work with the wavefunction in the non-relativistic limit. Taking the ansatz
@@
where $\xi_{n\ell m}=\omega_{n\ell m}-\mu$, $N_{n\ell m}$ is the occupation number of particle in the eigenstate, and $u_{n\ell m}({\bf r})$ is normalized to 1. To the leading order in $\alpha$, $u_{n\ell m}({\bf r})$ is given by the normalized hydrogenic eigen-wavefunction with a Bohr radius of $r_{B}\equiv GM/\alpha^{2}$.
The Klein-Gordon equation of $\psi$ also allows unbound states, which satisfy the out-going boundary condition at spatial infinity. The unbound states are continuous in spectrum, with the eigenfunctions, in the Newtonian limit, given by the stationary Coulomb waves [63]
@@
Figure 2: Parameter space of two non-interacting fields. Without loss of generality, we assume $\psi_{211}$ to be the fastest growing mode, which excludes the white regime. Representative examples of the evolution in some regions are shown in Fig. 3, while the boundaries and the superradiant evolution of each region are described in Sec. II.2. In particular, we find that superradiant growth of $\psi_{211}$ might be affected by $\varphi$, as which may accelerate black hole spin down, causing $\psi_{211}$ depletes prematurely.
Figure 3: Presentative examples of non-interacting two-field superradiant evolution. The upper panel shows the evolution of occupation numbers in the region of corresponding color in Fig. 2, while the lower panel shows the evolution of black hole spin. In the left example, $\varphi_{211}$ and $\varphi_{322}$ cannot grow sufficiently become the black hole spin drops quickly due to the superradiant growth of $\psi_{211}$. In the middle example, $\psi_{211}$, $\varphi_{211}$, $\psi_{322}$ and $\varphi_{322}$ undergo efficient superradiant growth one after another and dominate the black hole spin in sequence. In the right example, the superradiant growth of $\varphi_{211}$ is suppressed by that of $\psi_{211}$, while $\varphi_{322}$ and $\psi_{322}$ can still grow in order after $\psi_{211}$ saturates.
@@ II.2 Superradiance Evolution
where we have defined $\gamma_{i}=\Gamma_{i}/\mu$ and $\epsilon_{i}=N_{i}/GM^{2}$, the over-dot denotes derivative with respect to the dimensionless time $\tilde{t}\equiv\mu t$, and the summation over all bound states of both $\psi$ and $\varphi$.
In the case of single filed $\psi$, $\psi_{211}$ typically grows most rapidly with
@@
In the case of two fields, the superradiance evolution depends on the parameters. In Fig. 2, we show the parameter space of different types of evolution, assuming $\psi_{211}$ to be the fastest growing mode. [^2] The case that $\varphi_{211}$ grows faster can be obtained by replacing $q$ with $1/q$. Specially, if $\gamma_{\psi_{322}}>\gamma_{\varphi_{211}}$ and $\gamma_{\psi_{322}}>\gamma_{\varphi_{322}}$ (region I.A), $\psi_{322}$ will undergo efficient superradiant growth after $\psi_{211}$ saturates and dominate black hole spin down. This process is then followed by the superradiant growth of $\varphi_{211}$ and $\varphi_{322}$. If $\gamma_{\varphi_{211}}<\gamma_{\psi_{322}}<\gamma_{\varphi_{322}}$ and $\tilde{a}^{\varphi_{322}}_{\rm crit}>\tilde{a}^{\psi_{211}}_{\rm crit}$ (region I.B), $\varphi_{211}$ and $\varphi_{322}$ cannot grow sufficiently as the black hole spin drops below their critical spin quickly as $\psi_{211}$ grows, while $\psi_{322}$ can still grow sufficiently after $\psi_{211}$ saturates. If $\gamma_{\varphi_{211}}>\gamma_{\psi_{322}},\gamma_{\varphi_{322}}$ and $\tilde{a}^{\varphi_{211}}_{\rm crit}<\tilde{a}^{\psi_{211}}_{\rm crit}$ (region II), $\varphi_{211}$ will undergo efficient superradiant growth after $\psi_{211}$ saturates and dominate black hole spin down. This process is then followed by the superradiant growth of $\psi_{322}$ and $\varphi_{322}$. If $\gamma_{\varphi_{322}}>\gamma_{\psi_{322}}$ and $\tilde{a}^{\varphi_{211}}_{\rm crit}>\tilde{a}^{\psi_{211}}_{\rm crit},\tilde{a}^{\varphi_{322}}_{\rm crit}<\tilde{a}^{\psi_{211}}_{\rm crit}$ (region III), $\varphi_{211}$ will not grow, while $\varphi_{322}$ will grow efficiently after $\psi_{211}$ saturates, followed by the superradiant growth of $\psi_{322}$. A similar analysis can be extended in region IV, where higher modes of $\varphi$, such as $\varphi_{433}$, should also be taken into account. A representative example of each type is shown in Fig. 3. These examples are obtained by solving Eq. (10) together with $\dot{\epsilon}_{i}=\gamma_{i}\epsilon_{i}$, assuming the initial conditions of $\epsilon_{i}=10^{-10}$ when the black hole spin $\tilde{a}_{0}=0.99$. For simple, we treated the black hole mass (and hence $\alpha$) to be constant, given the fact that the black hole mass typically varies by a small amount during superradiant evolution. We find that, in region I and region IV, $\psi_{211}$ finally depletes after reaching the maximum occupation number, due to the development of $\varphi$ field.
@@ III Cubic Interactions
In this section, we investigate the superradiant growth of interacting scalar fields. We shall consider cubic interactions, which might take effects first comparing to higher dimensional interactions. We shall focus on the interactions between the fields instead of self-interactions, as which have been investigated in Ref. [8]. This leaves two relevant operators: $\psi^{2}\,\varphi$ and $\psi\,\varphi^{2}$. As the growth rate is extremely sensitivity to the mass of the fields, one of the fields usually grow much faster than another field, even if the masses of the two field are comparable. Without loss of generality, let $\psi$ be the one that grows faster, in which case $\psi^{2}\,\varphi$ is expected to dominate over $\psi\,\varphi^{2}$. For these reasons, we shall neglect $\psi\,\varphi^{2}$, and consider Lagrangian (1) with ${\cal L}_{\rm int}=\lambda\psi^{2}\,\varphi$, where $\lambda$ is the coupling constant.
@@
Assuming both fields are initially in vacuum states with small quantum fluctuations, the superradiance process can be investigated perturbatively: At the very beginning, interactions between the field is suppressed due to the low occupation numbers, and the superradiance process is the same as it in the case of free field. As superradiance continues, the occupation number in the fastest growing mode, i.e., ${\bar{\psi}}_{211}$, increases, resulting in a notable interaction between the two fields. Depending on the parameters, the interaction may lead to different processes: The $\psi$ field may lose energy to the black hole via the bound state of $\varphi$. It may also lose energy to the spatial infinity via the unbound states of $\varphi$ and $\psi$. The effects on the superradiant growth of such processes can be investigated by their corrections on the eigenfrequncies, which will be discussed as follows.
@@ III.1 Bound state interactions
We shall start with interactions between $\varphi$ and the bound states of $\psi$. It is convenient to described the bound states with $\bar{\psi}$} defined in the non-relativistic ansatz (6), and work in the weak gravity limit. By substituting ansatz (6) and expanding terms in powers of $\alpha$, Eqs. (13) and (14) can be written into the form of
@@
where $\bar{\lambda}\equiv\lambda/2\mu$ is the reduced coupling strength, and $\hat{H}\simeq-\tfrac{\nabla^{2}}{2\mu}-\tfrac{\alpha}{r}$ to the leading order of $\alpha$ (see App. A for the detailed derivation). Here we do not take the non-relativistic ansatz for $\varphi$, as its relativistic states could also contribute in this process.
To compute the corrections on the eigenfrequency, we take one of the $\bar{\psi}$} in the source terms as background ${\bar{\psi}}_{0}$. In our case, the background is mostly in the ${\bar{\psi}}_{211}$ mode, i.e., ${\bar{\psi}}_{0}=\sqrt{N_{0}}u_{211}(\mathbf{r})e^{-i\xi_{211}t}$. While ${\bar{\psi}}_{211}$ should in principle grow with time, given the fact that ${\rm Im}\xi_{211}\ll{\rm Re}\xi_{211}$, one can neglect the imaginary part of $\xi_{211}$ and treat ${\bar{\psi}}_{0}$ to be stationary in our perturbative calculation for simplicity [8]. It is convenient to use Feynman-like diagrams to represent the channels involved in Eqs. (LABEL:eq:eomm). In particular, $\bar{\psi}$} on the r.h.s. of Eqs. (LABEL:eq:eomm) corresponds to a leg on the l.h.s. of the vertex, and vice versa; and $\bar{\psi}$$^{*}$ on the r.h.s. of Eqs. (LABEL:eq:eomm) corresponds to a leg on the r.h.s. of the vertex, and vice versa. We are interested in the interaction corrections on the fastest growing mode ${\bar{\psi}}_{211}$. Given Eqs. (LABEL:eq:eomm), there are three possible channels, which are shown in Fig. 4.
@@
Figure 4: Possible channels involved in Eqs. (LABEL:eq:eomm). The left diagram corresponds to the s-channel described by Eqs. (16) and (17), while the middle and right diagrams correspond to the t- and u-channels described by Eqs. (32).
Figure 5: Numerical results of integrals (28), (30), and (42). In practice, we have to impose a cut-off $n_{\rm max}$ for the summation in Eq. (27). Given the left plot, we expect the summation converges well when $n_{\rm max}>8$ for $q<5$.
For the s-channel, we have
@@
We shall start with Eq. (17). To the leading order in $\bar{\lambda}$, the r.h.s. of Eq. (17) is given by $\bar{\lambda}{\bar{\psi}}^{(0)}{\bar{\psi}}_{0}e^{-2i\mu t}$. Then by considering the ansatz
@@
with $\hat{M}_{\sigma}$ can be obtained by replacing $\partial_{t}$ in $\Box$ with $i\sigma$ and $\tilde{s}\left(\mathbf{r}\right)=\bar{\lambda}\sqrt{N_{0}}u_{211}^{2}$. The above equation can be solved with the Green’s function method, [^3] Technically, the integral should be performed with the tortoise coordinates $\mathbf{r}_{*}$. See App. B for the precise formula. Nevertheless, we shall not distinguish $\mathbf{r}_{*}$ with $\mathbf{r}$ here, as $\tilde{s}(\mathbf{r})$ only has support around $r\sim GM/\alpha^{2}$ where $\mathbf{r}_{*}\simeq\mathbf{r}$.
@@
In App. B, we show that the above integral can be written as
@@
are the bound and scattering state wavefunctions of the free $\varphi$ field. Here the bound states has ${\rm Re}\sigma_{n\ell m}<\nu$ and are discrete in spectrum, while the scattering states has ${\rm Re}\sigma_{k\ell m}\geq\nu$ and are continuous in spectrum. [^4] Also see Ref. [64] for a recent discussion on the complete set of mode functions. The precise formulae for the coefficients $b_{n\ell m}$ and $b(k)_{\ell m}$ are derived in App. B. In practice, when we compute the numerical value of the coefficients, we approximate the wavefunctions $v_{n\ell m}$ and $v_{k\ell m}$ with their analytical expressions in the weak gravity limit (discussed in App. A), since $\tilde{s}(\mathbf{r})$ only has support far from the black hole horizon. It turns out that the coefficients are non-trivial only if $\ell=m=2$ due to the angular dependence, and are given by
@@
where $u_{n\ell m}$ denotes the bound states of $\bar{\psi}$}, and the summation excludes $u_{211}$ as which can be absorbed into ${\bar{\psi}}^{(0)}$. Substituting Eq. (25) into Eq. (16), and using the orthogonality of $u_{n\ell m}$, we can find $a_{n\ell m}=0$ and the frequency correction from the s-channel $\delta\omega_{211}^{s}$. Specifically, the correction can be written as $\delta\omega_{211}^{s}=\delta\omega_{211}^{sB}+\delta\omega_{211}^{sE}$ with
@@
which indicates that the superradiant growth rate of ${\bar{\psi}}_{211}$ is suppressed if it interacts with a decaying ${\bar{\varphi}}_{n22}$, as ${\bar{\varphi}}_{n22}$ may lose energy to the black hole in the s-channel, and vice versa. In the last line of Eq. (27), we considered that $\omega_{0}\simeq\mu$ and ${\rm Re}\sigma_{n}\simeq q\mu$, and have defined
@@
which is a dimensionless factor depending only on the mass ratio $q$, with its values shown in Fig. 5. For the corrections from unbound states, we find that $\delta\omega_{211}^{sE}$ has a non-zero imaginary part only if $k_{0}$ is real due to the optical theorem. Specially, by performing the integral over $k$ with residue theorem, one has
@@
which represents the superradiant growth of ${\bar{\psi}}_{211}$ is suppressed by their annihilation into radiation of $\varphi$, a process happening only if $2\omega_{0}\geq\nu$. Here
@@
and can be investigated in a similar way. The frequency corrections, however, happen to be real, and therefore will not be discussed in details.

### v3

# Black hole superradiance of interacting multifield

@@ Abstract
We investigate black hole superradiance evolution of the interacting multiple fields. We consider a model of two scalar fields interacting with a cubic coupling, and study the superradiant evolution of the cloud. We demonstrate that superradiance is typically suppressed when the superradiant field couples to another field, even with a very weak coupling strength. This implies that the constraints on dark particles derived from single-field analyses can be revised in the presence of interactions. Moreover, we find that the multifield superradiant evolution and its corresponding observational signatures can be different across parameter spaces, which makes black hole superradiance an even more powerful probe of the dark sector in particle physics.
@@ I Introduction
While the dark sector may consist of multiple particle species with nontrivial interactions, most studies assume the presence of only a single superradiant field. Although some works have considered interactions between the superradiant field and other fields, their focus has largely been on phenomena, such as photon polarization by the cloud [26, 25, 24] and annihilation of the cloud into photons [51], fermions [27] and scalars [41], with the assumption of a preformed superradiant cloud. Not only has the superradiance of the coupled fields (or generally, interactions between the black hole and bound states of the coupled fields) been overlooked, but its consequential effects on the superradiance of the primary field also lack investigation. Therefore, a systematic investigation of superradiant evolution of the interacting multiple fields is essential and necessary.
In this work, we investigate black hole superradiance in the presence of multiple fields. We consider a model of two scalar fields interacting with a cubic coupling, and study the superradiance evolution of the cloud. We demonstrate that superradiance is typically suppressed if the superradiant field couples to another field even with a very weak coupling strength. This implies that the constraints on dark particles derived from single-field analyses can be significantly revised in the presence of interactions. Moreover, the superradiant evolution in different regions of the parameter space are distinct. This makes black hole superradiance an even more powerful probe of the dark sector in particle physics.
The rest of this paper is organized as follows. We start with the superradiance evolution of two noninteracting fields in Sec. II. Then we shall consider a cubic interaction, and shall discuss its effects on the superradiant growth rate in Sec. III and the superradiant evolution in Sec. IV. Finally, we shall discussion the implications observations in Sec. V. Sec. VI is devoted to the conclusion and discussion. We will take (,+,+,+)(-,+,+,+) metric convention and set $\hbar=c=1$.
II Superradiance of free multifields
In this section, we investigate the superradiance process of noninteracting multifields (also see Ref. [44] for a similar study). For the purpose of demonstration, we shall start with a model of two massive scalar fields in the Kerr background, the Lagrangian of which is given by
@@ II.1 Eigenstates of free fields
where $x^{\mu}=(t,r,\theta,\varphi)$ are the Boyer-Lindquist coordinates, $R_{\omega\ell m}(r)$ is the radial function, and $S_{\omega\ell m}(\theta)$ turns out to be the spheroidal harmonics. The boundary conditions at the black hole horizon and at spatial infinity single out a set of eigenfrequencies. For bound states, i.e., states vanishes at spatial infinity, the eigenfrequencies $\omega$ are shown to be a set of discrete complex number $\omega_{n\ell m}$, labeled by three quantum numbers $n$, $\ell$, and $m$. These eigenfrequencies can be obtained numerically. For instance, Fig. 1 shows the eigenfrequencies of some bound states obtained with the continued fraction method [39]. For $\alpha\ll 1$, we have [13]
@@
Figure 1: Eigenfrequencies of bound states with $n\ell m$ being $211$ (in black) and $322$ (in red). The upper and lower plots show the real and imaginary parts of the eigenfrequencies respectively. The solid lines are given by Eqs. (3) and (4), while the dashed lines show the results obtained from numerical calculation.
While the wave function of the bound states can also be obtained numerically [40], for $\alpha\ll 1$, it is convenient to work with the wave function in the nonrelativistic limit. Taking the ansatz
@@
where $\xi_{n\ell m}=\omega_{n\ell m}-\mu$, $N_{n\ell m}$ is the occupation number of the particle in the eigenstate, and $u_{n\ell m}({\bf r})$ is normalized to 1. To the leading order in $\alpha$, $u_{n\ell m}({\bf r})$ is given by the normalized hydrogenic eigen-wave function with a Bohr radius of $r_{B}\equiv GM/\alpha^{2}$.
The Klein-Gordon equation of $\psi$ also allows unbound states, which satisfy the outgoing boundary condition at spatial infinity. The unbound states are continuous in spectrum, with the eigenfunctions, in the Newtonian limit, given by the stationary Coulomb waves [38]
@@
Figure 2: Parameter space of two noninteracting fields. Without loss of generality, we assume $\psi_{211}$ to be the fastest growing mode, which excludes the white regime. Representative examples of the evolution in some regions are shown in Fig. 3, while the boundaries and the superradiant evolution of each region are described in Sec. II.2. In particular, we find that superradiant growth of $\psi_{211}$ might be affected by $\varphi$, which may accelerate black hole spin down, causing $\psi_{211}$ depletes prematurely.
Figure 3: Presentative examples of noninteracting two-field superradiant evolution. The upper panel shows the evolution of occupation numbers in the region of corresponding color in Fig. 2, while the lower panel shows the evolution of black hole spin. In the left example, $\varphi_{211}$ and $\varphi_{322}$ cannot grow sufficiently because the black hole spin drops quickly due to the superradiant growth of $\psi_{211}$. In the middle example, $\psi_{211}$, $\varphi_{211}$, $\psi_{322}$ and $\varphi_{322}$ undergo efficient superradiant growth one after another and dominate the black hole spin in sequence. In the right example, the superradiant growth of $\varphi_{211}$ is suppressed by that of $\psi_{211}$, while $\varphi_{322}$ and $\psi_{322}$ can still grow in order after $\psi_{211}$ saturates.
@@ II.2 Superradiance evolution
where we have defined $\gamma_{i}=\Gamma_{i}/\mu$ and $\epsilon_{i}=N_{i}/GM^{2}$, the overdot denotes derivative with respect to the dimensionless time $\tilde{t}\equiv\mu t$, and the summation over all bound states of both $\psi$ and $\varphi$.
In the case of a single filed $\psi$, $\psi_{211}$ typically grows most rapidly with
@@
In the case of two fields, the superradiance evolution depends on the parameters. In Fig. 2, we show the parameter space of different types of evolution, assuming $\psi_{211}$ to be the fastest growing mode. [^2] The case that $\varphi_{211}$ grows faster can be obtained by replacing $q$ with $1/q$. Specially, if $\gamma_{\psi_{322}}>\gamma_{\varphi_{211}}$ and $\gamma_{\psi_{322}}>\gamma_{\varphi_{322}}$ (region I.A), $\psi_{322}$ will undergo efficient superradiant growth after $\psi_{211}$ saturates and dominates black hole spin down. This process is then followed by the superradiant growth of $\varphi_{211}$ and $\varphi_{322}$. If $\gamma_{\varphi_{211}}<\gamma_{\psi_{322}}<\gamma_{\varphi_{322}}$ and $\tilde{a}^{\varphi_{322}}_{\rm crit}>\tilde{a}^{\psi_{211}}_{\rm crit}$ (region I.B), $\varphi_{211}$ and $\varphi_{322}$ cannot grow sufficiently as the black hole spin drops below their critical spin quickly as $\psi_{211}$ grows, while $\psi_{322}$ can still grow sufficiently after $\psi_{211}$ saturates. If $\gamma_{\varphi_{211}}>\gamma_{\psi_{322}},\gamma_{\varphi_{322}}$ and $\tilde{a}^{\varphi_{211}}_{\rm crit}<\tilde{a}^{\psi_{211}}_{\rm crit}$ (region II), $\varphi_{211}$ will undergo efficient superradiant growth after $\psi_{211}$ saturates and dominates black hole spin down. This process is then followed by the superradiant growth of $\psi_{322}$ and $\varphi_{322}$. If $\gamma_{\varphi_{322}}>\gamma_{\psi_{322}}$ and $\tilde{a}^{\varphi_{211}}_{\rm crit}>\tilde{a}^{\psi_{211}}_{\rm crit},\tilde{a}^{\varphi_{322}}_{\rm crit}<\tilde{a}^{\psi_{211}}_{\rm crit}$ (region III), $\varphi_{211}$ will not grow, while $\varphi_{322}$ will grow efficiently after $\psi_{211}$ saturates, followed by the superradiant growth of $\psi_{322}$. A similar analysis can be extended in region IV, where higher modes of $\varphi$, such as $\varphi_{433}$, should also be taken into account. A representative example of each type is shown in Fig. 3. These examples are obtained by solving Eq. (10) together with $\dot{\epsilon}_{i}=\gamma_{i}\epsilon_{i}$, assuming the initial conditions of $\epsilon_{i}=10^{-10}$ when the black hole spin $\tilde{a}_{0}=0.99$. For simple, we treated the black hole mass (and hence $\alpha$) to be constant, given the fact that the black hole mass typically varies by a small amount during superradiant evolution. We find that, in region I and region IV, $\psi_{211}$ finally depletes after reaching the maximum occupation number, due to the development of the $\varphi$ field.
@@ III Cubic Interactions
In this section, we investigate the superradiant growth of interacting scalar fields. We shall consider cubic interactions which typically dominate over other higher dimensional interactions. We shall focus on the interactions between the fields instead of self-interactions, as those have been investigated in Ref. [9]. This leaves two relevant operators: $\psi^{2}\,\varphi$ and $\psi\,\varphi^{2}$. As the growth rate is extremely sensitive to the mass of the fields, one of the fields usually grow much faster than another field, even if the masses of the two field are comparable. Without loss of generality, let $\psi$ be the one that grows faster, in which case $\psi^{2}\,\varphi$ is expected to dominate over $\psi\,\varphi^{2}$. For these reasons, we shall neglect $\psi\,\varphi^{2}$, and consider Lagrangian (1) with ${\cal L}_{\rm int}=\lambda\psi^{2}\,\varphi$, where $\lambda$ is the coupling constant.
@@
Assuming both fields are initially in vacuum states with small quantum fluctuations, the superradiance process can be investigated perturbatively: At the very beginning, interactions between the field are suppressed due to the low occupation numbers, and the superradiance process is the same as it in the case of free field. As superradiance continues, the occupation number in the fastest growing mode, i.e., ${\bar{\psi}}_{211}$, increases, resulting in a notable interaction between the two fields. Depending on the parameters, the interaction may lead to different processes: The $\psi$ field may lose energy to the black hole via the bound state of $\varphi$. It may also lose energy to the spatial infinity via the unbound states of $\varphi$ and $\psi$. The effects on the superradiant growth of such processes can be investigated by their corrections on the eigenfrequencies, which will be discussed in what follows.
@@ III.1 Bound state interactions
We shall start with interactions between $\varphi$ and the bound states of $\psi$. It is convenient to described the bound states with $\bar{\psi}$} defined in the nonrelativistic ansatz (6), and work in the weak gravity limit. By substituting ansatz (6) and expanding terms in powers of $\alpha$, Eqs. (13) and (14) can be written into the form of
@@
where $\bar{\lambda}\equiv\lambda/2\mu$ is the reduced coupling strength, and $\hat{H}\simeq-\tfrac{\nabla^{2}}{2\mu}-\tfrac{\alpha}{r}$ to the leading order of $\alpha$ (see Appendix. A for the detailed derivation). Here we do not take the nonrelativistic ansatz for $\varphi$, as its relativistic states could also contribute in this process.
To compute the corrections on the eigenfrequency, we take one of the $\bar{\psi}$} in the source terms as background ${\bar{\psi}}_{0}$. In our case, the background is mostly in the ${\bar{\psi}}_{211}$ mode, i.e., ${\bar{\psi}}_{0}=\sqrt{N_{0}}u_{211}(\mathbf{r})e^{-i\xi_{211}t}$. While ${\bar{\psi}}_{211}$ should, in principle, grow with time, given the fact that ${\rm Im}\xi_{211}\ll{\rm Re}\xi_{211}$, one can neglect the imaginary part of $\xi_{211}$ and treat ${\bar{\psi}}_{0}$ to be stationary in our perturbative calculation for simplicity [9]. It is convenient to use Feynman-like diagrams to represent the channels involved in Eq. (LABEL:eq:eomm). In particular, $\bar{\psi}$} on the rhs of Eq. (LABEL:eq:eomm) corresponds to a leg on the lhs of the vertex, and vice versa; and $\bar{\psi}$$^{*}$ on the rhs of Eq. (LABEL:eq:eomm) corresponds to a leg on the rhs of the vertex, and vice versa. We are interested in the interaction corrections on the fastest growing mode ${\bar{\psi}}_{211}$. Given Eq. (LABEL:eq:eomm), there are three possible channels, which are shown in Fig. 4.
@@
Figure 4: Possible channels involved in Eq. (LABEL:eq:eomm). The left diagram corresponds to the $s$ channel described by Eqs. (16) and (17), while the middle and right diagrams correspond to the $t$ and $u$ channels described by Eq. (32).
Figure 5: Numerical results of integrals (28), (30), and (42). In practice, we have to impose a cutoff $n_{\rm max}$ for the summation in Eq. (27). Given the left plot, we expect the summation converges well when $n_{\rm max}>8$ for $q<5$.
For the $s$ channel, we have
@@
We shall start with Eq. (17). To the leading order in $\bar{\lambda}$, the rhs of Eq. (17) is given by $\bar{\lambda}{\bar{\psi}}^{(0)}{\bar{\psi}}_{0}e^{-2i\mu t}$. Then by considering the ansatz
@@
with $\hat{M}_{\sigma}$ can be obtained by replacing $\partial_{t}$ in $\Box$ with $i\sigma$ and $\tilde{s}\left(\mathbf{r}\right)=\bar{\lambda}\sqrt{N_{0}}u_{211}^{2}$. The above equation can be solved with the Green’s function method, [^3] Technically, the integral should be performed with the tortoise coordinates $\mathbf{r}_{*}$. See Appendix. B for the precise formula. Nevertheless, we shall not distinguish $\mathbf{r}_{*}$ with $\mathbf{r}$ here, as $\tilde{s}(\mathbf{r})$ only has support around $r\sim GM/\alpha^{2}$ where $\mathbf{r}_{*}\simeq\mathbf{r}$.
@@
In Appendix. B, we show that the above integral can be written as
@@
are the bound and scattering state wave functions of the free $\varphi$ field. Here the bound states have ${\rm Re}\sigma_{n\ell m}<\nu$ and are discrete in spectrum, while the scattering states have ${\rm Re}\sigma_{k\ell m}\geq\nu$ and are continuous in spectrum. [^4] Also see Ref. [32] for a recent discussion on the complete set of mode functions. The precise formulas for the coefficients $b_{n\ell m}$ and $b(k)_{\ell m}$ are derived in Appendix. B. In practice, when we compute the numerical value of the coefficients, we approximate the wave functions $v_{n\ell m}$ and $v_{k\ell m}$ with their analytical expressions in the weak gravity limit (discussed in Appendix. A), since $\tilde{s}(\mathbf{r})$ only has support far from the black hole horizon. It turns out that the coefficients are nontrivial only if $\ell=m=2$ due to the angular dependence, and are given by
@@
where $u_{n\ell m}$ denotes the bound states of $\bar{\psi}$}, and the summation excludes $u_{211}$ as which can be absorbed into ${\bar{\psi}}^{(0)}$. Substituting Eq. (25) into Eq. (16), and using the orthogonality of $u_{n\ell m}$, we can find $a_{n\ell m}=0$ and the frequency correction from the $s$ channel $\delta\omega_{211}^{s}$. Specifically, the correction can be written as $\delta\omega_{211}^{s}=\delta\omega_{211}^{sB}+\delta\omega_{211}^{sE}$ with
@@
which indicates that the superradiant growth rate of ${\bar{\psi}}_{211}$ is suppressed if it interacts with a decaying ${\bar{\varphi}}_{n22}$, as ${\bar{\varphi}}_{n22}$ may lose energy to the black hole in the $s$ channel, and vice versa. In the last line of Eq. (27), we considered that $\omega_{0}\simeq\mu$ and ${\rm Re}\sigma_{n}\simeq q\mu$, and have defined
@@
which is a dimensionless factor depending only on the mass ratio $q$, with its values shown in Fig. 5. For the corrections from unbound states, we find that $\delta\omega_{211}^{sE}$ has a nonzero imaginary part only if $k_{0}$ is real due to the optical theorem. Especially, by performing the integral over $k$ with residue theorem, one has
@@
which represents the superradiant growth of ${\bar{\psi}}_{211}$ suppressed by their annihilation into radiation of $\varphi$, a process happening only if $2\omega_{0}\geq\nu$. Here
@@
and can be investigated in a similar way. The frequency corrections, however, happen to be real, and therefore will not be discussed in detail.